OpenStudy (istim):

For an outdoor concert, a ticket price of $30 typically attracts 5000 people. For each $1 increase in the ticket price, 100 fewer people will attend. The revenue, R, is the product of the # of people attending and the price per ticket. If the ticket price that maximizes the revenue is $40, will that change if the concert area holds a max. of 1200 people? Explain.

OpenStudy (istim):

Formula for revenue is R=(5000-100x)(30+x), or R=-100x^2+2000+150 000. Derivative is -200x+2000. 2nd derivative is R"=-200. Max revenue is $160 000. Ticket price to Max. revenue is $40.

OpenStudy (apoorvk):

Why that second derivative? You just need to find the the 'x' for the maxima of the revenue function, which is basically equating the first derivative to zero.

OpenStudy (istim):

Just giving as much information as possible.

OpenStudy (apoorvk):

and remember 'x' is the *change* in price of each ticket, you still have to add/subtract it to the $30 to get the actual price of the ticket.

OpenStudy (istim):

x=10.

OpenStudy (apoorvk):

Yes! So ticket prices for optimum revenue for a typical sale of about 5000 tickets per day is 40 bucks. Now, if the constraint is that, the area itself can't hold more than 1200 people, then I am losing about 3800 people. so I can basically raise the price of each ticket by 38 bucks (100 people for every buck). So, find the out the revenue now, when x=38.

OpenStudy (istim):

How do I know that I am losing about 3800 people?

OpenStudy (apoorvk):

Well, the typical attendance (in the first line of the question) is about 5000. Now when the capacity is reduced to 1200 people, then you lose 5000-1200=3800 people.

OpenStudy (istim):

Oh. That logic surpasses me. Give me some time to try to work this out. Thanks for the help.

OpenStudy (apoorvk):

No worries, take it easy ;)

OpenStudy (istim):

Wait, how do I show this algebraically?

OpenStudy (apoorvk):

show algebraically what?

OpenStudy (istim):

I just got back to this question after goofing around, and I understand up to the 3 800 people part.

OpenStudy (apoorvk):

very well, your goofing around helped you i see. :p yes so, now 38000 aren't there. so for that amount of reduce in the audience, by how much can i afford to raise the ticket prices? it's written that for every 1 buck increase in the ticket price, i lose 1000 people. so if already lost 3800 people, how much can i raise the ticket price by? (because 1200 people are going to attend the play anyway)

OpenStudy (apoorvk):

actually anyway means that till i have an acceptable increase in tkt prices.

OpenStudy (istim):

I'm not really sure. So I just tried subbing 3800 as My quantity and sub it into Q=5000-100x, and then solve for x. Then I sub that into x for my price, and solve, were I get $68 instead.

OpenStudy (apoorvk):

nah!!!! 'x' is the decrease/increase in price per ticket. if i lose 1000 people for every dollar i increase, and i already lost 3800 people, then i can afford to increase the price by 38 dollars per ticket. now profit= (30+x)(5000-100x) 'x' here is 38. find out the max profit!

OpenStudy (apoorvk):

(which occurs for x=38)

OpenStudy (istim):

How did I figure out x?

OpenStudy (apoorvk):

i found out 'x' for you and showed you!! okay do you agree, that the '100x' in that equation is the decrease in the no. of attendees?

OpenStudy (apoorvk):

i mean attenders**

OpenStudy (istim):

Ok...

OpenStudy (apoorvk):

so get this?

OpenStudy (apoorvk):

100x=3800 = people lost due to the limited capacity of the audience. so, x = 3800/100 = 38!!

OpenStudy (istim):

Still wondering how you got =38...

OpenStudy (istim):

Where that equation come from? Did we change 5000 to 3800?

OpenStudy (apoorvk):

no!!!! 1200 is th limiting capacity of the auditorium, remember? so, 5000-1200 = 3800 people who we lost in the audience. we can'rt sell tickets to them, so we make up by increasing the prices to such a point that all tickets are sold, as well we can maximise our profit.

OpenStudy (istim):

Maybe we should start from the top again.

OpenStudy (apoorvk):

Sure!! Your job to read it all again. -.-

OpenStudy (apoorvk):

so, scroll up.

OpenStudy (istim):

That's a rather rude way to put it, but I'll re-read it.

OpenStudy (apoorvk):

Lol, didn't mean to be rude, sorry.

OpenStudy (istim):

It's alright. I'm just still confused on what exactly my next step should be.

OpenStudy (istim):

Cause I just stated my work a bit, and then we affirmed how the value 3800 was obtained. And then 100x=3800 popped up.

OpenStudy (apoorvk):

"For each $1 increase in the ticket price, 100 fewer people will attend." so, when 3800 fewer people are attending, what can the increase in the ticket price be?

OpenStudy (istim):

Am...I subbing in 3800 as the 'quantity' then?

OpenStudy (apoorvk):

no! in the original equation for profit, Profit = (30+x)(5000-100x) here 100x=3800 and x=38 so put in x=38 and get the max profit for this case.

OpenStudy (istim):

Like into the original equation right?

OpenStudy (apoorvk):

yes!

OpenStudy (istim):

And from that, my final answer right?

OpenStudy (istim):

AFK.

OpenStudy (apoorvk):

yes! what is the answer that you're getting?

OpenStudy (istim):

Answer for Price is $68. Revenue is $81 600

OpenStudy (apoorvk):

exactly! 30 + the increase, i.e. 38 is $68. substituting 38 in the original equation for x, gives you $81,600, which is correct as well.

OpenStudy (istim):

Ok. Thank you very much.

OpenStudy (apoorvk):

no worries!